The first week in July, the daily cost of electricity to a household was $4.24, $5.03, $5.23 and $3.99 for Monday through Thursday respectively. To have a five day average cost of no more than $5.00, what must Friday's cost be?
Answer by Theo(13342)

the 4 day cost was 4.24, 5.03, 5.23, 3.99.
the total for the 4 days was 18.49.
let x be the fifth cost.
the 5 day average would be equal to (18.49 + x) / 5.
if you want the 5 day average to be less than or equal to 5.00, then the formula to use would be:
5.00 >= (18.49 + x) / 5
this formula could also be shown as:
(18.49 + x) / 5 <= 5.00
multiply both sides of this equation by 5 to get:
18.49 + x <= 25.00
subtract 18.49 from both sides of this equation to get:
x <= 6.51
when x = 6.51, the 5 day average becomes (18.49 + 6.51) / 5 = 5.00
when x is less than 6.51, the 5 day average will be less than 5.00.
for example, when x = 6.00, the 5 day average becomes (18.49 + 6.00) / 5 = 4.898 which is less than 5.00.
your solution is that, to have a five day average cost of no more than $5.00, friday's cost must be smaller than or equal to $6.51
note that no more than means the same as less than or equal.